# Minimal Subsampled Rank-1 Lattices for Multivariate Approximation with Optimal Convergence Rate

### v2

@@ 4 Error bounds for subsampled rank-1 lattices
$S_{\mathcal{B}}^{\bm{X}_{J}}f\coloneqq\operatorname*{arg\,min}_{g\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}}\sum_{k\in J}\Big{|}g(\tfrac{k}{n}\bm{z}\bmod{}1)-f(\tfrac{k}{n}\bm{z}\bmod{}1)\Big{|}^{2}\,.$ (4.1)
@@
Throughout this paper we will use multisets for $J$, meaning that elements may repeat. For the cardinality $|J|$, elements are counted according to their repeated appearance, e.g. |$|\{1,1\}|=2$. Note that for the actual number of points used, or equivalently, the number of function evaluations used, it is not necessary to count by repeated appearance but it always provides an upper bound.
@@
$\bm{L}_{J,\mathcal{B}}\coloneqq\Big{[}\exp(2\pi\mathrm{i}\tfrac{k}{n}\langle\bm{h},\bm{z}\rangle)\Big{]}_{k\in J,\,\bm{h}\in\mathcal{B}}\in\mathds{C}^{|J|\times|\mathcal{B}|}$ (4.2)
@@
Let $\mathcal{B}\subseteq\mathds{Z}^{d}$ be a nonempty frequency index set for which the lattice $\bm{X}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ has the reconstructing property. For $J\subseteq\{0,\dots,n-1\}$ a multiset with $|J|\geq|\mathcal{B}|$, suppose $\bm{L}_{J,\mathcal{B}}$ from (4.2) has full rank $|\mathcal{B}|$. Then the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ defined in (4.1), with the subsampled lattice $\bm{X}_{J}$ is a projection onto $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$.
@@
By definition we have that $S_{\mathcal{B}}^{\bm{X}_{J}}$ maps linearly into $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$. To show the idempotence, we use the Fourier representation of a function supported on $\mathcal{B}$
$f=\sum_{\bm{h}\in\mathcal{B}}\widehat{f}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$
and collect the Fourier coefficients in the vector $\bm{\widehat{f}}_{\mathcal{B}}=[\widehat{f}_{\bm{h}}]_{\bm{h}\in\mathcal{B}}$. From the reconstructing property, there holds $\bm{f}_{J}=\bm{L}_{J,\mathcal{B}}\,\bm{\widehat{f}}_{\mathcal{B}}$. By Lemma 4.1 we have for the Fourier coefficients of the least squares approximation
@@
Before commenting on a specific choice in $J$, we relate the worst-case error of the least squares approximation to the spectral properties of certain matrices.
@@
$\bm{\Phi}_{J,\mathcal{B}}\coloneqq\Big{[}\frac{\exp(2\pi\mathrm{i}\frac{k}{n}\langle\bm{h},\bm{z}\rangle)}{\sqrt{r(\bm{h})}}\Big{]}_{k\in J,\,\bm{h}\notin\mathcal{B}}\,.$ (4.3)
@@
$\|S_{\mathcal{B}}^{\bm{X}_{J}}f-f\|_{L_{2}}^{2}=\|f-P_{\mathcal{B}}f\|_{L_{2}}^{2}+\|P_{\mathcal{B}}f-S_{\mathcal{B}}^{\bm{X}_{J}}f\|_{L_{2}}^{2}\,,$ (4.4)
@@
Now we estimate the second summand of (4.4). By Lemma 4.2, $S_{\mathcal{B}}^{\bm{X}_{J}}$ is a projection onto $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$ and we obtain using Lemma 4.1
@@
We have seen in Lemma 4.2 that it is possible to reconstruct trigonometric polynomials from values at a subsampled lattice $\bm{X}_{J}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k\in J}\subseteq\mathds{T}^{d}$, $J\subseteq\{0,\dots,n-1\}$ given that $\bm{L}_{J,\mathcal{B}}$ has full rank. The following lemma quantifies this for a randomly selected subset $J$.
@@
$\frac{|J|}{2}\leq\sigma_{\min}^{2}(\bm{L}_{J,\mathcal{B}})\quad\text{and}\quad\sigma_{\max}^{2}(\bm{L}_{J,\mathcal{B}})\leq\frac{3|J|}{2}\,.$ (4.6)
@@
$\|\bm{\Phi}_{J,\mathcal{B}}\|_{2}^{2}\leq 21\,(\log|J|+t)\sum_{\bm{h}\notin\mathcal{B}}\frac{1}{r(\bm{h})}+2\frac{|J|}{n}\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}\,.$ (4.7)
@@
$\frac{1}{2}\|f\|_{L_{2}}^{2}\leq\frac{1}{|J|}\sum_{k\in J}|f(\tfrac{k}{n}\bm{z}\bmod{}1)|^{2}\leq\frac{3}{2}\|f\|_{L_{2}}^{2}\quad\text{for all}\quad f\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}\,,$
where from their proof each bound holds with probability $1-\exp(-t)$, respectively. This random technique uses a matrix Chernoff concentration inequality [Tropp11, Theorem 1.1] with one of the first applications in the sampling context in [CM17, Theorem 2.1]. Using [BKPU22, Lemma 2.2] again converts this to the desired bounds on the singular values in (4.6).
@@
$\frac{6\,(\log|J|+t)}{|J|}\leq\frac{1}{|\mathcal{B}|}\leq\frac{13\,(\log|J|+t)}{|J|}\,.$ (4.8)
@@
We use the bound of the smallest singular value of $\bm{L}_{J,\mathcal{B}}$ and the bound on the spectral norm of $\bm{\Phi}_{J,\mathcal{B}}$ of Lemma 4.4. By Boole’s inequality, the overall probability exceeds $1-\exp(-t)-2^{3/4}\exp(-t)\geq 1-3\exp(-t)$, which is the claimed probability. Plugging in the spectral bounds into Theorem 4.3, we obtain
@@
[…]
@@
In Theorem 4.5 the spectral norm of an infinite-dimensional matrix $\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}$ occurs. We have
@@
where the first matrix product is infinite-dimensional, whereas the second matrix-product is finite-dimensional (may be used for computing the spectral norm numerically). In the following lemma we use the former to show that this quantity can be upper bounded in terms of the known quantity $\mathcal{S}_{n}(\bm{z})$ from (2.5). Note that the reconstructing property is not needed in this lemma.
@@
Now we fix the frequency set $\mathcal{B}$ to be as in Corollary 3.3, where we have guaranteed reconstructing property, allowing us to omit it in the assumptions.
@@
By Corollary 3.3 we know that $\bm{X}$ has the reconstructing property for $\mathcal{B}$. Thus, we can apply Theorem 4.5, where we plug in Lemma 4.6 for estimating $\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}$ and obtain with the desired probability
@@
By the definition of $\mathcal{B}$, we have $\sup_{\bm{h}\notin\mathcal{B}}1/r(\bm{h})\leq 2\sqrt{\mathcal{S}_{n}(\bm{z})}$, which, when plugged in the above estimate, yields the assertion. ∎
@@
Following the same proof, it is easily seen, that Theorem 4.7 still holds if we replace $\mathcal{S}_{n}(\bm{z})$ by any upper estimate, even though it changes the definition of $\mathcal{B}$ and with that the underlying method.
Note that the occurrence of the quantity $1/|\mathcal{B}|\sum_{\bm{h}\notin\mathcal{B}}1/r(\bm{h})$ in Theorem 4.7 is natural, since there exist function spaces where this is a lower bound on the sampling numbers with $|\mathcal{B}|$ samples, cf. [HKNV22, Theorem 2]. Comparing the full lattice algorithm Theorem 2.1 with Theorem 4.7, both bounds depend on the term $\sqrt{\mathcal{S}_{n}(\bm{z})}$. The difference is in the number of used samples being $|J|$ in the new algorithm instead of $n$, with $|J|$ usually far less than $n$, as we will see in Section 5.
5 Application to weighted Korobov spaces
@@
$H_{d,\alpha,\bm{\gamma}}\coloneqq\Big{\{}f\colon\mathds{T}^{d}\to\mathds{C}:\|f\|_{d,\alpha,\bm{\gamma}}^{2}=\sum_{\bm{h}\in\mathds{Z}^{d}}r_{d,\alpha,\bm{\gamma}}(\bm{h})\,|\widehat{f}_{\bm{h}}|^{2}<\infty\Big{\}}\,,$ (5.1)
@@
$r_{d,\alpha,\bm{\gamma}}(\bm{h})\coloneqq\gamma_{\operatorname{supp}(\bm{h})}^{-1}\prod_{j\in\operatorname{supp}(\bm{h})}|h_{j}|^{2\alpha}\quad\text{and}\quad\operatorname{supp}(\bm{h})\coloneqq\Big{\{}j\in\{1,\dots,d\}:h_{j}\neq 0\Big{\}}\,.$ (5.2)
@@
  • further including the smoothness parameter yields smoothness driven product and order dependent (SPOD) weights $\gamma_{\mathfrak{u}}=\sum_{\bm{\nu}_{\mathfrak{u}}\in\{1,\dots,\alpha\}^{|\mathfrak{u}|}}\Gamma_{\|\bm{\nu}_{\mathfrak{u}}\|_{1}}\prod_{j\in\mathfrak{u}}\gamma_{j,\nu_{j}}$, see e.g., [DKLNS14, Dick2016, Kaarnioja2020].
@@
It was already shown in [BKPU22], that the optimal polynomial rate can be recovered using a subset of the initial lattice given that the initial lattice is large enough. In particular, in [BKPU22] a lattice of size $n\gtrsim|J|^{2/(1-1/(2\alpha))}$ is necessary. We will show in Theorem 5.3 that the optimal polynomial rate $(|J|/\log|J|)^{-\alpha+\varepsilon}$ can be recovered using only an initial lattice of size $n\lesssim\,(|J|/\log|J|)^{2/\sqrt{1-\varepsilon/\alpha}}$ for any $0<\varepsilon<\alpha$.
In order to apply our general result on subsampled lattices, we first present known results for the full lattice. Note that the parameter $\alpha$ in [KMN24] is $2\alpha$ in this paper.
@@
$C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}\coloneqq\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\max\{1,|\mathfrak{u}|\}\,\gamma_{\mathfrak{u}}^{\lambda}\,(2\zeta(2\alpha\lambda))^{|\mathfrak{u}|}\,.$ (5.3)
@@
Applying the estimate on $\mathcal{S}_{n}(\bm{z})$ from Theorem 5.1 to Theorem 2.1 achieves the order of convergence $n^{-\alpha/2+\varepsilon}$ for the full lattice, which is half the optimal order.
@@
Then we have for $11/(2\alpha)<\lambda<$ 1 and $C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}$ from (5.3) that
@@
Note, in the original [KMN24, Lemma 5.2] the constant $C_{\lambda}$ equals $\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\gamma_{\mathfrak{u}}^{\lambda}\,(2\zeta(2\alpha\lambda))^{|\mathfrak{u}|}$. Here we used an upper estimate (5.3) to simplify our presentation.
@@
$e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})^{2}\leq 7\Big{(}1+\frac{C_{\lambda}^{1/\lambda}}{\gamma_{\{1\}}^{1/(2\alpha\lambda)}}\cdot\frac{\lambda}{1-\lambda}\Big{)}\,(13\,C_{\lambda})^{\frac{1}{2\alpha\lambda^{2}}}\,\Big{(}\frac{\log|J|+t}{|J|}\Big{)}^{\frac{1}{2\alpha\lambda^{2}}}$ (5.4)
for all $11/(2\alpha)<\lambda<$ 1, with $C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}$ from (5.3). Moreover, if $\bm{z}$ is constructed according to [KMN24, Algorithm 3.2], we have (with $\varphi(n)$ again denoting the Euler totient function)
$(\varphi(n))^{\frac{1}{4\alpha\lambda}}\,\log\varphi(n)\lesssim|J|\lesssim n^{\alpha\lambda}\log n\,,$ (5.5)
@@
$\Big{(}\frac{|J|}{\log|J|}\Big{)}^{\frac{1}{\alpha\lambda}}\lesssim n\quad\mbox{and}\quad\varphi(n)\lesssim\Big{(}\frac{|J|}{\log|J|}\Big{)}^{4\alpha\lambda}$ (5.6)
@@
For the index set ${\mathcal{B}}$ with radius $M:=1/(2\sqrt{\mathcal{S}_{n}({\bm{z}})})\geq 1$, Lemma 5.2 gives for all $11/(2\alpha)<\lambda<$ 1,
@@
Theorem 4.7 and Lemma 5.2 then yield
@@
Next we relate $|{\mathcal{B}}|$ to the bounds on $M=1/(2\sqrt{\mathcal{S}_{n}(\bm{z})})$ from Theorem 5.1, to obtain
$\Big{(}\frac{\varphi(n)}{\tau_{\lambda}\,C_{\lambda}^{2}}\Big{)}^{\frac{1}{4\alpha\lambda}}\Big{(}\frac{\gamma_{\{1\}}}{2}\Big{)}^{\frac{1}{2\alpha}}\leq\Big{(}\frac{\gamma_{\{1\}}}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big{)}^{\frac{1}{2\alpha}}\leq|\mathcal{B}|\leq C_{\lambda}\Big{(}\frac{1}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big{)}^{\lambda}\leq\frac{C_{\lambda}\,n^{\alpha\lambda}}{(8\zeta(2\alpha)\gamma_{\{1\}})^{\lambda/2}}\,.$
@@
In particular, for $n$ prime, we have $(|J|/\log|J|)^{1/(\alpha\lambda)}\lesssim n\lesssim(|J|/\log|J|)^{4\alpha\lambda}$. Substituting $1/(4\alpha\lambda^{2})=\alpha-\varepsilon$ for $0<\varepsilon<\alpha$, we obtain the convergence rate $(|J|/\log|J|)^{-\alpha+\varepsilon}$ with $(|J|/\log|J|)^{2\sqrt{1-\varepsilon/\alpha}}\lesssim n\lesssim(|J|/\log|J|)^{2/\sqrt{1-\varepsilon/\alpha}}$, where the exponent is arbitrarily close to $2$.

### v3

@@ 4. Error bounds for subsampled rank-1 lattices
(4.1) $S_{\mathcal{B}}^{\bm{X}_{J}}f\coloneqq\operatorname*{arg\,min}_{g\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}}\sum_{k\in J}\Big|g(\tfrac{k}{n}\bm{z}\bmod{}1)-f(\tfrac{k}{n}\bm{z}\bmod{}1)\Big|^{2}\,.$
@@
Throughout this paper we will use multisets for $J$, meaning that elements may repeat. For the cardinality $|J|$, elements are counted according to their repeated appearance, e.g. $|\{1,1\}|=2$. Note that for the actual number of points used, or equivalently, the number of function evaluations used, it is not necessary to count by repeated appearance but doing so always provides an upper bound.
@@
(4.2) $\bm{L}_{J,\mathcal{B}}\coloneqq\Big[\exp(2\pi\mathrm{i}\tfrac{k}{n}\langle\bm{h},\bm{z}\rangle)\Big]_{k\in J,\,\bm{h}\in\mathcal{B}}\in\mathds{C}^{|J|\times|\mathcal{B}|}$
@@
Let $\mathcal{B}\subseteq\mathds{Z}^{d}$ be a nonempty frequency index set for which the lattice $\bm{X}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k=0}^{n-1}\subseteq\mathds{T}^{d}$ has the reconstructing property. For $J\subseteq\{0,\dots,n-1\}$ a multiset with $|J|\geq|\mathcal{B}|$, suppose $\bm{L}_{J,\mathcal{B}}$ from (4.2) has full rank $|\mathcal{B}|$. Then the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ defined in (4.1), with the subsampled lattice $\bm{X}_{J}$, is a projection onto $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$.
@@
By definition we have that $S_{\mathcal{B}}^{\bm{X}_{J}}$ maps linearly into $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$. To show the idempotence, we use the Fourier representation of a function supported on $\mathcal{B}$, i.e.,
$f=\sum_{\bm{h}\in\mathcal{B}}\widehat{f}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}\,$,
and collect the Fourier coefficients in the vector $\bm{\widehat{f}}_{\mathcal{B}}=[\widehat{f}_{\bm{h}}]_{\bm{h}\in\mathcal{B}}$. From the reconstructing property, there holds $\bm{f}_{J}=\bm{L}_{J,\mathcal{B}}\,\bm{\widehat{f}}_{\mathcal{B}}$. By Lemma˜4.1 we have for the Fourier coefficients of the least squares approximation
@@
Before commenting on a specific choice of $J$, we relate the worst-case error of the least squares approximation to the spectral properties of certain matrices.
@@
(4.3) $\bm{\Phi}_{J,\mathcal{B}}\coloneqq\Big[\frac{\exp(2\pi\mathrm{i}\frac{k}{n}\langle\bm{h},\bm{z}\rangle)}{\sqrt{r(\bm{h})}}\Big]_{k\in J,\,\bm{h}\notin\mathcal{B}}\,.$
@@
(4.4) $\|S_{\mathcal{B}}^{\bm{X}_{J}}f-f\|_{L_{2}}^{2}=\|f-P_{\mathcal{B}}f\|_{L_{2}}^{2}+\|P_{\mathcal{B}}f-S_{\mathcal{B}}^{\bm{X}_{J}}f\|_{L_{2}}^{2}\,.$
@@
Now we estimate the second summand of (4.4). By Lemma˜4.2, $S_{\mathcal{B}}^{\bm{X}_{J}}$ is a projection onto $\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}$ and we obtain using Lemma˜4.1
@@
We have seen in Lemma˜4.2 that it is possible to reconstruct trigonometric polynomials from values at a subsampled lattice $\bm{X}_{J}=\{\frac{k}{n}\bm{z}\bmod{}1\}_{k\in J}\subseteq\mathds{T}^{d}$ for $J\subseteq\{0,\dots,n-1\}$ given that $\bm{L}_{J,\mathcal{B}}$ has full rank. The following lemma quantifies this for a randomly selected subset $J$.
@@
(4.6) $\frac{|J|}{2}\leq\sigma_{\min}^{2}(\bm{L}_{J,\mathcal{B}})\quad\text{and}\quad\sigma_{\max}^{2}(\bm{L}_{J,\mathcal{B}})\leq\frac{3|J|}{2}\,.$
@@
(4.7) $\|\bm{\Phi}_{J,\mathcal{B}}\|_{2}^{2}\leq 21\,(\log|J|+t)\sum_{\bm{h}\notin\mathcal{B}}\frac{1}{r(\bm{h})}+2\frac{|J|}{n}\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}\,.$
@@
$\frac{1}{2}\|f\|_{L_{2}}^{2}\leq\frac{1}{|J|}\sum_{k\in J}|f(\tfrac{k}{n}\bm{z}\bmod{}1)|^{2}\leq\frac{3}{2}\|f\|_{L_{2}}^{2}\;\;\text{for all}\;\;f\in\operatorname{span}\{\exp(2\pi\mathrm{i}\langle\bm{h},\cdot\rangle)\}_{\bm{h}\in\mathcal{B}}\,$,
where from their proof each bound holds with probability $1-\exp(-t)$, respectively. This random technique uses a matrix Chernoff concentration inequality [Tropp11, Theorem 1.1] with one of the first applications in the sampling context appearing in [CM17, Theorem 2.1]. Using [BKPU22, Lemma 2.2] again converts this to the desired bounds on the singular values in (4.6).
@@
(4.8) $\frac{6\,(\log|J|+t)}{|J|}\leq\frac{1}{|\mathcal{B}|}\leq\frac{13\,(\log|J|+t)}{|J|}\,.$
@@
We use the bound of the smallest singular value of $\bm{L}_{J,\mathcal{B}}$ and the bound on the spectral norm of $\bm{\Phi}_{J,\mathcal{B}}$ of Lemma˜4.4. By Boole’s inequality, the overall probability exceeds $1-\exp(-t)-2^{3/4}\exp(-t)\geq 1-3\exp(-t)$, which is the claimed probability. Plugging in the spectral bounds into Theorem˜4.3, we obtain
@@
We are ready to prove an upper bound on $(\log j+t)/j$ in terms of $1/b$ when $j:=\lceil\,12\,b\,(\log b+t)\,\rceil\geq 12\,b\,(\log b+t)$. Since $(\log j+t)/j$ is increasing with decreasing $j$, it suffices to show an estimate for $\ell\leq j$ such that
@@
In Theorem˜4.5 the spectral norm of an infinite-dimensional matrix $\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}$ occurs. We have
@@
where the first matrix product is infinite-dimensional, whereas the second matrix-product is finite-dimensional (and may be used for computing the spectral norm numerically). In the following lemma we use the former to show that this quantity can be upper bounded in terms of the known quantity $\mathcal{S}_{n}(\bm{z})$ from (2.5). Note that the reconstructing property is not needed in this lemma.
@@
Now we fix the frequency set $\mathcal{B}$ to be as in Corollary˜3.4, where we have guaranteed reconstructing property, allowing us to omit it in the assumptions.
@@
By Corollary˜3.4 we know that $\bm{X}$ has the reconstructing property for $\mathcal{B}$. Thus, we can apply Theorem˜4.5, where we plug in Lemma˜4.6 for estimating $\|\bm{\Phi}_{\mathcal{B}}\|_{2}^{2}$ and obtain with the desired probability
@@
By the definition of $\mathcal{B}$, we have $\sup_{\bm{h}\notin\mathcal{B}}(1/r(\bm{h}))\leq 2\sqrt{\mathcal{S}_{n}(\bm{z})}$, which, when plugged in the above estimate, yields the assertion. ∎
@@
Following the same proof, it is easily seen, that Theorem˜4.7 still holds if we replace $\mathcal{S}_{n}(\bm{z})$ by any upper estimate, even though it changes the definition of $\mathcal{B}$ and with that the underlying method.
Note that the occurrence of the quantity $(1/|\mathcal{B}|)\sum_{\bm{h}\notin\mathcal{B}}(1/r(\bm{h}))$ in Theorem˜4.7 is natural, since there exist function spaces where this is a lower bound on the sampling numbers with $|\mathcal{B}|$ samples, cf. [HKNV22, Theorem 2]. Comparing the full lattice algorithm Theorem˜2.1 with Theorem˜4.7, both bounds depend on the term $\sqrt{\mathcal{S}_{n}(\bm{z})}$. The difference is in the number of used samples being $|J|$ in the new algorithm instead of $n$, with $|J|$ usually far less than $n$, as we will see in Section˜5.
5. Application to weighted Korobov spaces
@@
(5.1) $H_{d,\alpha,\bm{\gamma}}\coloneqq\Big\{f\colon\mathds{T}^{d}\to\mathds{C}:\|f\|_{d,\alpha,\bm{\gamma}}^{2}=\sum_{\bm{h}\in\mathds{Z}^{d}}r_{d,\alpha,\bm{\gamma}}(\bm{h})\,|\widehat{f}_{\bm{h}}|^{2}<\infty\Big\}\,$,
@@
(5.2) $r_{d,\alpha,\bm{\gamma}}(\bm{h})\coloneqq\gamma_{\operatorname{supp}(\bm{h})}^{-1}\prod_{j\in\operatorname{supp}(\bm{h})}|h_{j}|^{2\alpha}\quad\text{and}\quad\operatorname{supp}(\bm{h})\coloneqq\Big\{j\in\{1,\dots,d\}:h_{j}\neq 0\Big\}\,.$
@@
  • including the smoothness parameter yields smoothness driven product and order dependent (SPOD) weights $\gamma_{\mathfrak{u}}=\sum_{\bm{\nu}_{\mathfrak{u}}\in\{1,\dots,\alpha\}^{|\mathfrak{u}|}}\Gamma_{\|\bm{\nu}_{\mathfrak{u}}\|_{1}}\prod_{j\in\mathfrak{u}}\gamma_{j,\nu_{j}}$, see e.g., [DKLNS14, Dick2016, Kaarnioja2020].
@@
It was already shown in [BKPU22], that the optimal polynomial rate can be recovered using a subset of the initial lattice given that the initial lattice is large enough. In particular, in [BKPU22] a lattice of size $n\gtrsim|J|^{2/(1-1/(2\alpha))}$ is necessary. We will show in Theorem˜5.3 and Corollary˜5.4 that the optimal polynomial rate $|J|^{-\alpha+\varepsilon}$ can be recovered using only an initial lattice of size $n\lesssim\,|J|^{2/\sqrt{1-\varepsilon/\alpha}}$ for any $0<\varepsilon<\alpha$.
In order to apply our general result on subsampled lattices, we first present known results for the full lattice. Note that the parameter $\alpha$ in [KMN24] is replaced by $2\alpha$ in this paper.
@@
(5.3) $C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}\coloneqq\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\max\{1,|\mathfrak{u}|\}\,\gamma_{\mathfrak{u}}^{\lambda}\,(2\zeta(2\alpha\lambda))^{|\mathfrak{u}|}\,.$
@@
Applying the estimate on $\mathcal{S}_{n}(\bm{z})$ from Theorem˜5.1 to Theorem˜2.1 achieves the order of convergence $n^{-\alpha/2+\varepsilon}$ for the full lattice, which is half the optimal order.
@@
Then we have for $1/(2\alpha)<\lambda<1$ and $C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}$ from (5.3) that
@@
In the original [KMN24, Lemma 5.2] the constant $C_{\lambda}$ equals $\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\gamma_{\mathfrak{u}}^{\lambda}\,(2\zeta(2\alpha\lambda))^{|\mathfrak{u}|}$. Here we used an upper estimate (5.3) to simplify our presentation.
@@
(5.4) $e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})^{2}\leq 7\Big(1+\frac{C_{\lambda}^{1/\lambda}}{\gamma_{\{1\}}^{1/(2\alpha\lambda)}}\cdot\frac{\lambda}{1-\lambda}\Big)\,(13\,C_{\lambda})^{\frac{1}{2\alpha\lambda^{2}}}\,\Big(\frac{\log|J|+t}{|J|}\Big)^{\frac{1}{2\alpha\lambda^{2}}}$
for all $1/(2\alpha)<\lambda<1$, with $C_{\lambda}=C_{d,\alpha,\bm{\gamma},\lambda}$ from (5.3). Moreover, if $\bm{z}$ is constructed according to [KMN24, Algorithm 3.2], we have (with $\varphi(n)$ again denoting the Euler totient function)
(5.5) $(\varphi(n))^{\frac{1}{4\alpha\lambda}}\,\log\varphi(n)\lesssim|J|\lesssim n^{\alpha\lambda}\log n\,$,
@@
(5.6) $\Big(\frac{|J|}{\log|J|}\Big)^{\frac{1}{\alpha\lambda}}\lesssim n\quad\mbox{and}\quad\varphi(n)\lesssim\Big(\frac{|J|}{\log|J|}\Big)^{4\alpha\lambda}\,.$
@@
For the index set $\mathcal{B}$} with radius $M:=1/(2\sqrt{\mathcal{S}_{n}({\bm{z}})})\geq 1$, Lemma˜5.2 gives for all $1/(2\alpha)<\lambda<1$,
@@
Theorem˜4.7 and Lemma˜5.2 then yield
@@
Next we relate $|{\mathcal{B}}|$ to the bounds on $M=1/(2\sqrt{\mathcal{S}_{n}(\bm{z})})$ from Theorem˜5.1, to obtain
$\Big(\frac{\varphi(n)}{\tau_{\lambda}\,C_{\lambda}^{2}}\Big)^{\frac{1}{4\alpha\lambda}}\Big(\frac{\gamma_{\{1\}}}{2}\Big)^{\frac{1}{2\alpha}}\leq\Big(\frac{\gamma_{\{1\}}}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big)^{\frac{1}{2\alpha}}\leq|\mathcal{B}|$ $\leq C_{\lambda}\Big(\frac{1}{2\sqrt{\mathcal{S}_{n}(\bm{z})}}\Big)^{\lambda}$
$\leq\frac{C_{\lambda}\,n^{\alpha\lambda}}{(8\zeta(2\alpha)\gamma_{\{1\}})^{\lambda/2}}\,.$